OpenStudy (anonymous):

PLEASE HELP!! 1 5/8 - 1 1/8=?

OpenStudy (anonymous):

Can you convert those to improper fractions?

OpenStudy (ghazi):

\[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\] use this

OpenStudy (anonymous):

13/8 - 9/8

OpenStudy (ghazi):

great

OpenStudy (anonymous):

now what?

OpenStudy (ghazi):

\[\frac{ 13 }{ 8 }-\frac{ 9 }{ 8 }=\frac{ 13-9 }{ 8 }\] do it

OpenStudy (anonymous):

4/8=1/2

OpenStudy (anonymous):

4/8?

OpenStudy (ghazi):

yes ....reduce it further

OpenStudy (anonymous):

its 1/2 right?

OpenStudy (anonymous):

lol yes its 1/2

OpenStudy (anonymous):

K :)

OpenStudy (anonymous):

9/16 + 3/4 =?

OpenStudy (anonymous):

sorry i cant do your home homework for you haha

OpenStudy (anonymous):

i just need help

OpenStudy (anonymous):

is it 3/4?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

use a calculator find out?

OpenStudy (anonymous):

thats impossible

OpenStudy (anonymous):

do what ghazi said ...and no it is possible lmao.

OpenStudy (anonymous):

asdfghjkl;